A callable bond has 15 years to maturity and can be called in 5 years. The bond’s coupon rate is 12% with semi-annual coupon payments. The par value is $1000. If the bond is called, the call price will be $1100. The bond is currently selling for $1055.35 What is the difference between its yield-to-call and yield-to-maturity?
Answer #1

Hello Sir/ Mam

Given that:

Periods to Maturity = 30

Periods to Maturity = 10

Coupon = $60

Future Maturity Value = $1,000

Call Price = $1,100

PV = $1,055.35

Now, using excel function, "=RATE(30,-60,1055.35,-1000,0)*2", we get YTM = 11.23%

Using excel function, "=RATE(10,-60,1055.35,-1100,0)*2", we get YTM = 12.01%

Hence, the difference between YTM and YTC = 0.78%

I hope this solves your doubt.

Feel free to comment if you still have any query or need something else. I'll help asap.

Do give a thumbs up if you find this helpful.